OpenStudy (anonymous):

For first order linear differential equation problems, where do I start? How do i use P(x) and Q(x) and u(x)???

ganeshie8 (ganeshie8):

maybe start wid an example problem :)

OpenStudy (anonymous):

I have a test wednesday morn..... still not understanding! and yes, I know about the IF method but not quite absorbing

OpenStudy (anonymous):

ok uhm (3x+sin2x)dx -dy=0

OpenStudy (anonymous):

And it just says to solve ^^

ganeshie8 (ganeshie8):

your goal is to find \(y\)

ganeshie8 (ganeshie8):

when it just says "solve"

ganeshie8 (ganeshie8):

\( (3x+\sin2x)dx -dy=0 \) \( (3x+\sin2x)dx = dy\) Integrate both sides \( \int (3x+\sin2x)dx = \int dy\) \( \int (3x+\sin2x)dx = y\)

ganeshie8 (ganeshie8):

we're done. we have "solved" y

OpenStudy (anonymous):

I know how to solve that way, but... is there a way to solve it with all the fancy P(x) and Q(x) and u(x) stuff? I think this may be slightly connected to the IF thing that we talked about yesterday before your lunch ^^

ganeshie8 (ganeshie8):

sure u can solve it by finding IF and multiplying it both sides but nobody does that cuz this can be easily solvable by separating variables

ganeshie8 (ganeshie8):

if u want to do it, we can do it... just to give u practice with IF method ;)

ganeshie8 (ganeshie8):

step1 : put the given equation into "standard linear differential equation form" : \(\large y' + p(x) y = q(x)\)

OpenStudy (anonymous):

hmmm what do I do if I come across something like dy/dx=3y+sin2x??? there's two terms, and I can't really transfer a term over because that would mean excluding it from the dx or dy that it's supposed to be with...

ganeshie8 (ganeshie8):

this is a new equation eh ?

ganeshie8 (ganeshie8):

Are we done with the first equation ?

OpenStudy (anonymous):

OoO no I just rearranged it in a failed attempt to make it into the linear form

ganeshie8 (ganeshie8):

nope, u changed the second equation. look closer

ganeshie8 (ganeshie8):

your original equation : (3x+sin2x)dx -dy=0 your rearranged equation : dy/dx=3y+sin2x

ganeshie8 (ganeshie8):

they dont match (you have replaced 3x with 3y)

ganeshie8 (ganeshie8):

which one is correct ?

OpenStudy (anonymous):

oh whoops!!! It's 3y~

ganeshie8 (ganeshie8):

cool :) then none of the discussion so far is meaningless

ganeshie8 (ganeshie8):

you cannot solve it by using "variable separation method"

ganeshie8 (ganeshie8):

you will hve to use the IF method

OpenStudy (anonymous):

XP srry haha yeah~

ganeshie8 (ganeshie8):

lets start from here : dy/dx=3y+sin2x

OpenStudy (anonymous):

oh hey! haha accidental success!

ganeshie8 (ganeshie8):

you wanto change it to form : \(y' + p(x) y = q(x) \)

ganeshie8 (ganeshie8):

:) replace dy/dx wid y' first

OpenStudy (anonymous):

so then just take the 3y and slap it over onto the left side?

ganeshie8 (ganeshie8):

\( \large \frac{dy}{dx}=3y+\sin2x \) \( \large y'=3y+\sin2x \) \( \large y'+ (-3)y = \sin2x \)

ganeshie8 (ganeshie8):

yes ! compare it wid the standard form : \(y' + p(x) y = q(x)\)

ganeshie8 (ganeshie8):

p(x) = ? q(x) = ?

OpenStudy (anonymous):

p(x) is -3, and q(x) is sin2x~

ganeshie8 (ganeshie8):

yup ! step1 : find the IF

ganeshie8 (ganeshie8):

step2 : multiply IF both sides and compress the left hand side

ganeshie8 (ganeshie8):

step3 : integrate both sides

ganeshie8 (ganeshie8):

for step1, use the formula : IF = \(\large e^{\int p dx}\)

OpenStudy (anonymous):

ohh ok

ganeshie8 (ganeshie8):

for our present problem, IF = \(e^{\int -3 dx} = e^{-3x}\)

ganeshie8 (ganeshie8):

you're done with step1

ganeshie8 (ganeshie8):

next, multiply the IF both sides of ur rearranged equation in standard form

OpenStudy (anonymous):

okay, so I got to the part where I've already condensed the left side and put the whole thing under an integration thing, and so now it looks like this: ye^-3x=(integral sign) e^-3x (sin2x) -----is there supposed to be a dx here???

ganeshie8 (ganeshie8):

\(\large y'+ (-3)y = \sin2x\) multiply IF = \(e^{-3x}\) both sides \(\large e^{-3x}y'+ e^{-3x}(-3)y = e^{-3x}\sin2x\) \(\large (ye^{-3x})' = e^{-3x}\sin2x\) Integrate both sides \(\large \int (ye^{-3x})'~ dx = \int e^{-3x}\sin2x~dx\)

OpenStudy (anonymous):

so the dx just kind of magically appears along with the integration signs?

OpenStudy (anonymous):

And I have no idea how to integrate the right side...

ganeshie8 (ganeshie8):

\(\int dx\) we're integrating both sides w.r.t x, so dx also comes

ganeshie8 (ganeshie8):

integrating right hand side is a different thing, but, u are comfortable wid the mechanics of IF method right ?

OpenStudy (anonymous):

so what about the left side? isn't it integration w.r.t. y?

ganeshie8 (ganeshie8):

nope, both sides we're integrating w.r.to x only

ganeshie8 (ganeshie8):

\(\large y'+ (-3)y = \sin2x\) multiply IF = \(e^{-3x}\) both sides \(\large e^{-3x}y'+ e^{-3x}(-3)y = e^{-3x}\sin2x\) \(\large (ye^{-3x})' = e^{-3x}\sin2x\) Integrate both sides \(\large \int (ye^{-3x})'~ dx = \int e^{-3x}\sin2x~dx\) \(\large ye^{-3x} = \int e^{-3x}\sin2x~dx\)

OpenStudy (anonymous):

ohhh ok, so it's much like putting a derivative onto the equation~

ganeshie8 (ganeshie8):

yes, lets look at integrating right hand side

ganeshie8 (ganeshie8):

u familiar wid "integration by parts" eh ?

OpenStudy (anonymous):

wait so would the derivative sign on the left side come off even w.r.t y?

ganeshie8 (ganeshie8):

we're interested in finding "y" as a function of x, so "integrating with respect to x" annihilates the "derivative with respect to x"

ganeshie8 (ganeshie8):

you always integrate with respect to the independent variable : x or t

OpenStudy (anonymous):

hmmm so then the left side isn't mulitplied by dy/dx?

ganeshie8 (ganeshie8):

\(\large \int (ye^{-3x})'~ dx = \int e^{-3x}\sin2x~dx\)

ganeshie8 (ganeshie8):

let me ask u a question :) wat does \(\large (ye^{-3x})'\) represent ?

OpenStudy (anonymous):

and so the left side is [ ye^-3e ]dy/dx? or no???

ganeshie8 (ganeshie8):

\(\large (ye^{-3x})'\) is same as \(\large \frac{d}{dx}(ye^{-3x})\)

ganeshie8 (ganeshie8):

just a notation thing, for me the first notation looks neat

ganeshie8 (ganeshie8):

so i use that, if u prefer the d/dx 's u may use that as well

OpenStudy (anonymous):

oh hmmm haha ^^ I thought the first thing meant dy/dx, so I was confused~

OpenStudy (anonymous):

ok so then the left side is left bare with nthn, and now on to the right side?

ganeshie8 (ganeshie8):

\(\large y'+ (-3)y = \sin2x\) multiply IF = \(e^{-3x}\) both sides \(\large e^{-3x}y'+ e^{-3x}(-3)y = e^{-3x}\sin2x\) \(\large (ye^{-3x})' = e^{-3x}\sin2x\) Integrate both sides \(\large \int (ye^{-3x})'~ dx = \int e^{-3x}\sin2x~dx\) \(\large ye^{-3x} = \int e^{-3x}\sin2x~dx \)

ganeshie8 (ganeshie8):

you are at peace wid everything so far eh ?

OpenStudy (anonymous):

yesh, more or less so ^^

ganeshie8 (ganeshie8):

the right side is just a "integration problem" : you can do it by parts very easily : http://www.wolframalpha.com/input/?i=%5Cint+e%5E%28-3x%29+sin%282x%29+dx

OpenStudy (anonymous):

WOAH COOL wat is thissss

ganeshie8 (ganeshie8):

\(\large y'+ (-3)y = \sin2x\) multiply IF = \(e^{-3x}\) both sides \(\large e^{-3x}y'+ e^{-3x}(-3)y = e^{-3x}\sin2x\) \(\large (ye^{-3x})' = e^{-3x}\sin2x\) Integrate both sides \(\large \int (ye^{-3x})'~ dx = \int e^{-3x}\sin2x~dx\) \(\large ye^{-3x} = \int e^{-3x}\sin2x~dx \) \(\large ye^{-3x} = \frac{-1}{13}e^{-3x}[3\sin(2x) + 2\cos(2x)] + c \)

ganeshie8 (ganeshie8):

we're done

OpenStudy (anonymous):

QQ I still don't understand how wolf ram got the answer tho...

ganeshie8 (ganeshie8):

dont wry, we're going to work it now

ganeshie8 (ganeshie8):

\(\int e^{-3x} \sin (2x) ~dx\)

ganeshie8 (ganeshie8):

familiar with "by parts" ?

OpenStudy (anonymous):

kind of~ but I only know the + ones~ where you split the terms?

ganeshie8 (ganeshie8):

\(\int uv' = uv - \int u'v \)

ganeshie8 (ganeshie8):

does that look familiar /

ganeshie8 (ganeshie8):

?

OpenStudy (anonymous):

can't say it does... |:

OpenStudy (anonymous):

maybe it's a BC topic...

OpenStudy (anonymous):

^^ well, thank you so much for your help~! I'll be going to bed now! it's 4am already... *yawn* bibi^^ have a good... lunchtime? LOL I meant afterenon ^^

ganeshie8 (ganeshie8):

\(\large I = \int e^{-3x} \sin (2x) ~dx \) By parts : \(u = \sin(2x), v' = e^{-3x}\) , gives \(\large I = \sin (2x) \frac{-e^{-3x} }{3} - \int 2\cos (2x) \frac{-e^{-3x} }{3} ~dx \) \(\large I = \sin (2x) \frac{-e^{-3x} }{3} +\frac{2}{3} \int \cos (2x) e^{-3x} ~dx \)

ganeshie8 (ganeshie8):

do it by parts again for that integral again

ganeshie8 (ganeshie8):

\(\large I = \sin (2x) \frac{-e^{-3x} }{3} +\frac{2}{3} \int \cos (2x) e^{-3x} ~dx \) \(\large I = \sin (2x) \frac{-e^{-3x} }{3} +\frac{2}{3} \left( \cos(2x) \frac{-e^{-3x} }{3} - \int (-2\sin (2x))\frac{-e^{-3x} }{3} \right) \) \(\large I = \sin (2x) \frac{-e^{-3x} }{3} -\frac{2}{9} \cos(2x)e^{-3x} - \frac{2}{3} \int \sin (2x)e^{-3x} \) \(\large I = \sin (2x) \frac{-e^{-3x} }{3} -\frac{2}{9} \cos(2x)e^{-3x} - \frac{2} {3} I \)

ganeshie8 (ganeshie8):

solve \(I\)

ganeshie8 (ganeshie8):

have good sleep :)

ganeshie8 (ganeshie8):

^^Thats one very long method :/

ganeshie8 (ganeshie8):

Another easy/super fast way(my fav) is to work out \(\int e^{-3x}\sin(2x)dx\) using complex numbers : \(\large \mathbb{\int e^{-3x} \sin(2x) dx = Img \left( \int e^{-3x} e^{i (2x)}dx\right) }\) \(\large \mathbb{~~~~~~~ = Img \left( \int e^{x(2i-3)} dx\right)}\) \(\large \mathbb{~~~~~~~ = Img \left( \frac{e^{x(2i-3)}}{2i-3} \right) }\) \(\large \mathbb{~~~~~~~ = Img \left( e^{-3x}\frac{2i+3}{13}[ e^{i(2x)}] \right) }\) \(\large \mathbb{~~~~~~~ = Img \left( e^{-3x}\frac{2i+3}{13}[ \cos(2x) + i\sin(2x)] \right) }\) \(\large \mathbb{~~~~~~~ = \frac{e^{-3x}}{13} [ 2\cos(2x) + 3\sin(2x)] }\)